Suppose that at a busy traffic junction, the probability p of an individual car having an accident is 0.0001. During a certain part of the day, 100 cars pass through the junction. What is the probability of two or more cars being involved in an accident within this period?
1
Expert's answer
2020-06-01T19:12:41-0400

λ=np=100*0.0001=0.01

P(x≥2)=1-P(x=0)-P(x=1)=1-e-0.01-0.01e-0.01=1-0.99-0.0099=0.0001